please help to solve equation
1/x+1/6=1/4

Answer by nerdybill(7384)

1/x+1/6=1/4
.
Multiply both sides by a common denominator 24x:
24x[1/x+1/6] = 24x[1/4]
24 + 4x = 6x
24 = 2x
12 = x